OpenStudy (anonymous):

1. y = x + 6 y = 2x (1 point) (6, 12) (–12, –6) (–6, –12) (2, 4)

OpenStudy (anonymous):

For questions 1–3, what is the solution of the system? Use substitution.

OpenStudy (anonymous):

x=6 y=12 1 is the ans

OpenStudy (anonymous):

2. 3x + 2y = 7 –3x + 11 = y (1 point)

OpenStudy (anonymous):

3. 2x – y = –7 4x – y = –4 (1 point) (–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2)

OpenStudy (anonymous):

4. –4y + 4 = x 2x + 8y = 8 (1 point) one solution two solutions infinitely many solutions no solution 5. 6x + 2 = y 3y – 18x = 12 (1 point) one solution two solutions infinitely many solutions no solution For questions 4–5, how many solutions does the system have?

OpenStudy (anonymous):

2) y=-4 x= 5

OpenStudy (anonymous):

3) b

OpenStudy (anonymous):

THANK YOU! YOU ARE A LIFESAVER!! :D

OpenStudy (anonymous):

4) infinite many solutions

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

5)one solution